OpenStudy (anonymous):

consider the following equation 7x+6y=17..find the equation in slope-intercept form, for the line which is parallel to this line and passes through the point (-6,2)

OpenStudy (anonymous):

parallel lines have the same slope. So our lie must have a slope of -7/6 Our line is, y-2=(-7/6)(x+6) rearrange to get in slope intercept form

OpenStudy (anonymous):

so y=(-7/6)x-5

OpenStudy (anonymous):

so the 6 in the slope and the point cancel out....then add 2 to the other side leaving -5?

OpenStudy (anonymous):

in the x portion of the point, yes :)

OpenStudy (anonymous):

yay!!! you all are awesome on here!!!!

OpenStudy (anonymous):

:)

OpenStudy (ricardovg):

Parallel = same slope 7x+6y=17 => y = (17-7x)/6 => y=(17/6)-(7/6)x then m is -(7/6) (y-2)/(x+6) = -(7/6) y=-(7/6)x-5 then new y pass by (-6,2) and has slope -(7/6)